# Gyroscopically stabilized quantum spin rotors

### v1

@@ Abstract
The spin probability rephases for $\tau=\pi/\omega_{\gamma}$, where $\lambda_{\tau}=1$, due to a perfect overlap of the states of the two branches, reduced only by NV spin decoherence with realistic dephasing time of $T_{2}=10\,\mu\mathrm{s}$ [58, 59, 60]. This can be seen in Fig. 3(b), which demonstrates that increasing the rotation frequency $\omega$ enhances the interference signal by reducing the total interferometer time. Fig. 3(c) shows the total interferometer time and the recurrence probability $P_{\uparrow}(\tau=\pi/\omega_{\gamma})$ as a function of the spin-rotor coupling $g/D_{\rm nv}$. The shaded region indicates where the NV’s $T_{2}$ time exceeds the interferometer time, resulting in a significant enhancement of the interference signal.
The proposed experiments for gyroscopically stabilized spin-rotors are within reach given the recent achievements in the field of levitated particles [45, 17, 19, 18, 31, 53]. Electrically charged nanodiamonds hosting NV centers have been trapped, manipulated, and monitored both via light scattering as well as by preparation and measurements of the embedded spins (state initialization with a green laser and microwave driving to the magnetic states) [44, 30, 61]. Gyroscopic spinning of charged particles with embedded spin degrees of freedom was recently demonstrated [31, 53], achieving MHz rotation rates. Illuminating the particle with circularly polarized light can further boost the rotation rate beyond GHz [62, 63, 64].
In summary, we discussed the gyroscopic stabilization of rapidly spinning nanodiamonds with a single embedded NV spin, leading to an effective spin-rotor Hamiltonian containing only a single rotational degree of freedom. We proposed several schemes to use this Hamiltonian to probe spin-rotation coupling in state-of-the-art setups and to perform mechanical quantum interference experiments. Our work presents the first steps towards applications of gyroscopically stabilized spin rotors.
Acknowledgments— We thank Cosimo C. Rusconi, Sanchar Sharma, Maxime Perdriat, Julen Pedernales, and Klaus Hornberger for stimulating discussions and comments on the manuscript. VW and BAS acknowledge funding by the Carl-Zeiss-Foundation through the project QPhoton. VW acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Training Group 1995 ‘Quantum Many-Body Methods in Condensed Matter Systems’ Project Number 240766775. VW and SVK are supported by the German Federal Ministry of Education and Research (BMBF) project QECHQS (Grant No. 16KIS1590K). BAS is supported by the DFG–510794108.
@@ Appendix A Rotation tensor and spin-rotor Hamiltonian
$[J_{\lambda},J_{\mu}]=-i\hbar\varepsilon_{\lambda\mu\nu}J_{\nu},\quad{\rm and}\quad[S_{\lambda},S_{\mu}]=i\hbar\varepsilon_{\lambda\mu\nu}S_{\nu},$ (18a)
@@
$[J_{k},J_{\ell}]=i\hbar\varepsilon_{k\ell m}J_{m}\quad{\rm and}\quad[S_{k},S_{\ell}]=i\hbar\varepsilon_{k\ell m}S_{m},$ (19a)
@@
$[J_{k},S_{\ell}]=i\hbar\varepsilon_{k\ell m}S_{m}.$ (19b)
Note that $[J_{j},J_{\mu}]=0$. Thus the operators $({\bf J}^{2},J_{3},J_{z},{\bf S}^{2},S_{\bf n})$ form a complete set of commuting observables [65].
@@
$p_{\alpha}=$ ${\bf J}\cdot{\bf e}_{z}$ (20a)
$p_{\beta}=$ $\frac{1}{2}{\bf J}\cdot\left(-\sin\alpha{\bf e}_{x}+\cos\alpha{\bf e}_{y}\right)+{\rm h.c.}$ (20b)
@@ Appendix B Trapping potential in the rotating frame
${\bf E}({\bf r},t)=\frac{U_{\rm ac}}{d_{0}}\cos(\omega_{\rm ac}t)\sum_{j=x,y,z}A_{j}({\bf e}_{j}\cdot{\bf r}){\bf e}_{j},$ (22)
@@
×$\times[{\bf n}_{\mu}(\Omega)\cdot{\bf e}_{j}][{\bf n}_{\nu}(\Omega)\cdot{\bf e}_{j}]\right]^{2}.$ (24)
@@
and Eq. (17) reduces to the Hamiltonian (Gyroscopically stabilized quantum spin rotors). We move to a frame co-rotating with the rotor $UHU^{\dagger}+i\hbar\partial_{t}UU^{\dagger}$ with the unitary operator
@@
which adds the term $-\omega p_{\alpha}$ and replaces $p_{\alpha}$ by $p_{\alpha}+I\omega$ while leaving the remaining Hamiltonian unchanged. Using that in the rotating frame $p_{\alpha}=0$, the Hamiltonian reads
@@
+$+\left\{\dfrac{I\omega-p_{\gamma}\cos\beta}{2I\sin^{2}\beta},S_{1}\cos\gamma-S_{2}\sin\gamma\right\}+V_{\rm qu}(\beta)$
@@
Here, a number of terms can be neglected: First, the terms $(S_{1}^{2}+S_{2}^{2})/2I$ and $S_{3}^{2}/2I_{3}$ are negligible in comparison to the NV zero-field splitting $D_{\rm nv}S_{1}^{2}/\hbar$ and can thus be dropped. Second, the spin-rotation coupling in $\beta$ as described by the last term in the first line is always suppressed compared to the $\beta$ trapping potential as well as to the gyroscopic potential in $\gamma$ and can thus be neglected. Third, the spin-rotation coupling in $\gamma$ as described by $-p_{\gamma}S_{3}/I_{3}$ as well as by the second line is always smaller than the gyroscopic coupling described by the second term in the first line and can thus also be dropped. The simplified Hamiltonian reads
@@ Appendix C Adiabatic approximation
Figure 4: Dynamics of $\gamma$ for the spin state $|S_{z}=-\hbar\rangle$ as a function of $\omega_{0}t$, comparing the solution of (a) the gyroscopic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (light blue, solid) with the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) and (b) the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) with the dispersive Hamiltonian (12) (light blue, solid). The nanodiamond has an ellipsoidal shape with semi-axes $l_{3}=200$ nm and $l_{1}=l_{2}=0.4l_{3}$, $B=-102\,\mathrm{mT}$, $\omega/2\pi=0.1\,\mathrm{MHz}$, $\omega_{0}=\omega_{\gamma}$ and $\sigma_{\gamma}=\sqrt{\hbar/2I_{\mathrm{eff}}\omega_{\gamma}}$. We approximated the Fock space as a finite space of dimension $d=40$ and chose for both the initial states of the $\beta$- and $\gamma$-oscillator a coherent state $|\alpha_{\beta,\gamma}=0.1\rangle$.
The fast $\xi$-librations in Eq. (Gyroscopically stabilized quantum spin rotors) can now be adiabatically eliminated by expanding an arbitrary spin-rotor state $|\phi\rangle$ in the oscillator eigenbasis $|n\rangle$ with $H_{\xi}|n\rangle=\hbar\omega(n+1/2)|n\rangle$ for $\omega_{\xi}\simeq\omega$. Using the ansatz (5),
@@
Here we used that $[p_{\gamma},W]=0$ and defined the coupling operators
@@
$+(\omega-\gamma_{0}B)\left[S_{1}\cos\left(\tilde{\gamma}\right)-S_{2}\sin\left(\tilde{\gamma}\right)\right].$ (35)
@@
By expressing the cosine and sine functions in terms of exponential functions and defining the overlap integrals (7), the effective Hamiltonian $H_{n}$ governing the dynamics of $|\chi_{n}\rangle$ can be written in the form (Gyroscopically stabilized quantum spin rotors). In Fig. 4(a) we compare the dynamics of $\gamma$ close to $\gamma=0$ for Eq. (Gyroscopically stabilized quantum spin rotors) and Eq. (Gyroscopically stabilized quantum spin rotors).
Appendix D Barnett anti-alignment

### v2

@@ Abstract
The spin probability rephases for $\tau=\pi/\omega_{\gamma}$, where $\lambda_{\tau}=1$, due to a perfect overlap of the states of the two branches, reduced only by NV spin decoherence with the realistic dephasing time of $T_{2}=10\,\mu\mathrm{s}$ [69, 70, 71]. Other environmental decoherence channels remain negligible in the regime considered (see App. I). This can be seen in Fig. 3(b), which demonstrates that increasing the rotation frequency $\omega$ enhances the interference signal by reducing the total interferometer time. Figure 3(c) shows the total interferometer time and the recurrence probability $P_{\uparrow}(\tau=\pi/\omega_{\gamma})$ as a function of the spin-rotor coupling $g/D_{\rm nv}$. The shaded region indicates where the NV’s $T_{2}$ time exceeds the interferometer time, resulting in a significant enhancement of the interference signal.
The proposed experiments for gyroscopically stabilized spin rotors are within reach given the recent achievements in the field of levitated particles [48, 17, 19, 18, 32, 64]. Electrically charged nanodiamonds hosting NV centers have been trapped, manipulated, and monitored both via light scattering as well as by preparation and measurements of the embedded spins (state initialization with a green laser and microwave driving to the magnetic states [72]) [47, 31, 73]. Gyroscopic spinning of charged particles with embedded spin degrees of freedom was recently demonstrated [32, 64], achieving MHz rotation rates. Illuminating the particle with circularly polarized light can further boost the rotation rate beyond GHz [74, 75, 76].
In summary, we discussed the gyroscopic stabilization of rapidly spinning nanodiamonds with a single embedded NV spin, leading to an effective spin-rotor Hamiltonian containing only a single rotational degree of freedom. We proposed several schemes to use this Hamiltonian to probe spin-rotation coupling in state-of-the-art setups and to perform mechanical quantum interference experiments. Our work presents the first steps toward applications of gyroscopically stabilized spin rotors.
Acknowledgments— We thank Cosimo C. Rusconi, Sanchar Sharma, Maxime Perdriat, Julen Pedernales, and Klaus Hornberger for stimulating discussions and comments on the manuscript. V.W. and B.A.S. acknowledge funding by the Carl-Zeiss-Foundation through the Project QPhoton. V.W. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Training Group 1995 ‘Quantum Many-Body Methods in Condensed Matter Systems’ Project No. 240766775. V.W. and S.V.K. are supported by the German Federal Ministry of Research, Technology and Space (BMFTR) Project QECHQS under Grant No. 16KIS1590K. B.A.S. is supported by the DFG–510794108.
@@ Appendix A Rotation tensor and spin-rotor Hamiltonian
$J_{\lambda},J_{\mu}]=-i\hbar\varepsilon_{\lambda\mu\nu}J_{\nu},\quad{\rm and}\quad[S_{\lambda},S_{\mu}]=i\hbar\varepsilon_{\lambda\mu\nu}S_{\nu},$ (18a)
@@
$J_{k},J_{\ell}]=i\hbar\varepsilon_{k\ell m}J_{m}\quad{\rm and}\quad[S_{k},S_{\ell}]=i\hbar\varepsilon_{k\ell m}S_{m},$ (19a)
@@
$J_{k},S_{\ell}]=i\hbar\varepsilon_{k\ell m}S_{m}.$ (19b)
Note that $J_{j},J_{\mu}]=0$. Thus the operators $({\bf J}^{2},J_{3},J_{z},{\bf S}^{2},S_{\bf n})$ form a complete set of commuting observables [77].
@@
$p_{\alpha}=$ ${\bf J}\cdot{\bf e}_{z},$ (20a)
$p_{\beta}=$ $\frac{1}{2}{\bf J}\cdot\left(-\sin\alpha{\bf e}_{x}+\cos\alpha{\bf e}_{y}\right)+{\rm h.c.},$ (20b)
@@
We emphasize that these canonical angular momentum operators differ from the purely kinetic angular momentum $\mathbf{L}$ that appears in the configuration-space description based on the Euler-angle velocities $\dot{\alpha},\dot{\beta},\dot{\gamma}$. The Hamiltonian used throughout the manuscript is formulated in terms of the canonical phase-space variables and therefore naturally incorporates the spin contribution to the total angular momentum $\mathbf{J}=\mathbf{L}+\mathbf{S}$.
@@ Appendix B Trapping potential in the rotating frame
${\bf E}({\bf r},t)=\frac{U_{\rm ac}}{d^{2}_{0}}\cos(\omega_{\rm ac}t)\sum_{j=x,y,z}A_{j}({\bf e}_{j}\cdot{\bf r}){\bf e}_{j},$ (22)
@@
[…]
@@
and Eq. (17) reduces to the Hamiltonian (Gyroscopically stabilized quantum spin rotors). For an electric trap with $U_{\mathrm{ac}}=2.5\,\mathrm{kV}$, $\omega_{\mathrm{ac}}/2\pi=0.5\,\mathrm{MHz}$, electrode spacing $d_{0}=350\,\mu\mathrm{m}$ [65], particle dimensions $l_{3}=200\,\mathrm{nm}$, $l_{1}=l_{2}=0.3l_{3}$, and surface charge density $\sigma=3.5\,\mu\mathrm{C}/\mathrm{m}^{2}$, the resulting harmonic frequency of the trapping potential in $\beta$ compared to a rotation frequency $\omega/2\pi=1\,\mathrm{MHz}$ is
$\omega_{\beta}=\sqrt{\frac{U_{\mathrm{ac}}^{2}(Q-Q_{3})^{2}}{8I^{2}\omega_{\mathrm{ac}}^{2}d_{0}^{4}}}\approx 4\times 10^{-5}\,\omega,$ (26)
hence $\omega_{\xi}\simeq\omega$. We move to a frame corotating with the rotor $UHU^{\dagger}+i\hbar\partial_{t}UU^{\dagger}$ with the unitary operator
@@
which adds the term $\omega p_{\alpha}$ and replaces $p_{\alpha}$ by $p_{\alpha}+I\omega$ while leaving the remaining Hamiltonian unchanged. Using that in the rotating frame $p_{\alpha}=0$, the Hamiltonian reads
@@
$+\left\{\dfrac{I\omega-p_{\gamma}\cos\beta}{2I\sin^{2}\beta},S_{1}\cos\gamma-S_{2}\sin\gamma\right\}+V_{\rm qu}(\beta)$
@@
Here, a number of terms can be neglected: First, the terms $(S_{1}^{2}+S_{2}^{2})/2I$ and $S_{3}^{2}/2I_{3}$ are negligible in comparison to the NV zero-field splitting $D_{\rm nv}S_{1}^{2}/\hbar$ and can thus be dropped. Second, the spin-rotation coupling in $\beta$ as described by the last term in the first line is always suppressed compared to the $\beta$ trapping potential as well as to the gyroscopic potential in $\gamma$ and can thus be neglected. Third, the spin-rotation coupling in $\gamma$ as described by $p_{\gamma}S_{3}/I_{3}$ as well as by the second line is always smaller than the gyroscopic coupling described by the second term in the first line. Considering $S_{3}=\hbar$ and small fluctuations in $\beta$ characterized by the zero-field amplitude $\xi_{0}=\sqrt{\hbar/2I\omega}$, the spin-rotation coupling compares to the gyroscopic coupling $p_{\gamma}\omega\xi$ as $S_{3}/I_{3}\omega\xi_{0}\approx 10^{-4}$ ($l_{3}=200\,\mathrm{nm}$, $l_{1}=l_{2}=0.3l_{3}$, $\omega/2\pi=1\,\mathrm{MHz}$), and thus can be dropped. The simplified Hamiltonian reads
@@ Appendix C Adiabatic approximation
Figure 4: Dynamics of $\gamma$ for the spin state $|S_{1}=$-$\hbar\rangle$ as a function of $\omega_{0}t$, comparing the solution of (a) the gyroscopic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (light blue, solid) with the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) and (b) the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) with the dispersive Hamiltonian (12) (light blue, solid). The nanodiamond has an ellipsoidal shape with semiaxes $l_{3}=200$ nm and $l_{1}=l_{2}=0.4l_{3}$, $B=-102\,\mathrm{mT}$, $\omega/2\pi=0.1\,\mathrm{MHz}$, and $\sigma_{\gamma}=\sqrt{\hbar/2I_{\mathrm{eff}}\omega_{\gamma}}$. We approximated the Fock space as a finite space of dimension $d=40$ and chose for both the initial states of the $\beta$- and $\gamma$-oscillator a coherent state $|\alpha_{\beta,\gamma}=0.1\rangle$.
The fast $\xi$ librations in Eq. (Gyroscopically stabilized quantum spin rotors) can now be adiabatically eliminated by expanding an arbitrary spin-rotor state $|\phi\rangle$ in the oscillator eigenbasis $|n\rangle$ with $H_{\xi}|n\rangle=\hbar\omega(n+1/2)|n\rangle$ for $\omega_{\xi}\simeq\omega$. Using the ansatz (5),
@@
Here we used that $p_{\gamma},W]=0$ and defined the coupling operators
@@
[…]
@@
By expressing the cosine and sine functions in terms of exponential functions and defining the overlap integrals (7), the effective Hamiltonian $H_{n}$ governing the dynamics of $|\chi_{n}\rangle$ can be written in the form (Gyroscopically stabilized quantum spin rotors), where we neglected the term $\propto S_{3}^{2}$. Using experimentally accessible conditions ($l_{3}=200\,\mathrm{nm}$, $l_{1}=l_{2}=0.3l_{3}$, $\omega/2\pi=1\,\mathrm{MHz}$, T=300T=300\,K, $B=100\,\mathrm{mT}$), we find $n=k_{\mathrm{B}}T/\hbar\omega\approx 6.25\times 10^{6}$ such that $I^{2}\omega^{4}/\hbar^{2}\gamma_{0}^{2}B^{2}\approx 3.8\times 10^{12}\gg n$, and likewise, the ratio $\hbar\gamma_{0}B/I\omega^{2}\approx 5\times 10^{-7}\ll 1$. Thus, the overlap integrals can be approximately set to unity ($f_{n},g_{n}\simeq 1$), for which the effective Hamiltonian (Gyroscopically stabilized quantum spin rotors) follows. In Fig. 4(a) we compare the dynamics of $\gamma$ close to $\gamma=0$ for Eq. (Gyroscopically stabilized quantum spin rotors) and Eq. (Gyroscopically stabilized quantum spin rotors).
C.1 Validity regime of the $\beta$-elimination
Figure 5: Validity of the adiabatic elimination of the $\xi$-mode. Parameter region satisfying Eq. (38) as a function of the rotation frequency $\omega$ and the major semiaxis length $l_{3}$, with $l_{1}=0.2l_{3}$, $B=-100\,\mathrm{mT}$ and different temperatures $T$ determining the mean occupation $n_{\gamma}$.
The adiabatic elimination of the fast $\beta$ libration mode requires a well-defined separation of timescales between the $\beta$ and $\gamma$ degrees of freedom. Quantitatively, the elimination requires
$\left|\dfrac{p_{\gamma}}{I\omega}\right|\ll 1,\quad\text{and}\quad\left|\dfrac{\hbar\gamma_{0}B}{I\omega^{2}}\right|\ll 1,$ (38)
where the first condition suppresses transitions between distinct $\beta$ levels and ensures that the $\beta$-mode follows the slow $\gamma$-dynamics adiabatically, while the second inequality guarantees that the overlap integrals Eq. (7) can be set to unity. Using $p_{\gamma}=\sqrt{\hbar n_{\gamma}I_{3}\omega_{\gamma}}$ with $n_{\gamma}$ the mean thermal occupation of $\gamma$ at temperature $T$, we quantitatively map the regime where both criteria in Eq. (38) are below 0.01. Fig. 5 shows the resulting parameter region as a function of the rotation frequency $\omega$ and particle size $l_{3}$ for $l_{1}=0.2l_{3}$ and $B=-100\,\text{mT}$. Larger aspect ratios or weaker fields further further improve the separation of scales.
C.2 Validity of neglecting the Zeeman-induced $S_{3}p_{\gamma}$ coupling
To quantify the impact of the Zeeman-induced spin-rotation coupling term $\gamma_{0}BS_{3}p_{\gamma}/I\omega$, we compare Fig. 6 the full coupled $\gamma$-spin dynamics including this term [Eq. (Gyroscopically stabilized quantum spin rotors)] with the resulting trajectories of those obtained from the simplified effective Hamiltonian in which it is omitted. The simulations were initialized in a spin superposition and evaluated for magnetic fields of 100100\,mT. For $\omega/2\pi\geq 1\,\mathrm{MHz}$ and $l_{3}=100\,$nm, the influence of the coupling term decreases rapidly, producing only minor quantitative deviations. For $l_{3}=200\,$nm the deviations already fall below numerical resolution. Thus, the Zeeman-induced coupling is strongly suppressed in the experimentally relevant regime.
[…]
Appendix D Imperfections of the Rotor–Spin Model
D.1 NV-axis misalignment
[…]
A small tilt of the NV quantization axis relative to the principal inertia axis modifies the zero-field splitting term of the Hamiltonian. Considering the NV axis as ${\bf n}={\bf n}_{1}+\varepsilon{\bf n}_{3}$ with $\varepsilon\ll 1$, the zero-field splitting becomes
$H_{\mathrm{zfs}}=\dfrac{D_{\mathrm{nv}}}{\hbar}(S_{1}+\varepsilon S_{3})^{2}\approx\dfrac{D_{\mathrm{nv}}}{\hbar}\left(S_{1}^{2}+\varepsilon\{S_{1},S_{3}\}\right),$ (39)
so that the effective Hamiltonian (Gyroscopically stabilized quantum spin rotors) acquires the perturbation
$\delta H_{\mathrm{eff}}$ $=\dfrac{\varepsilon D_{\mathrm{nv}}}{\hbar}W^{\dagger}\{S_{1},S_{3}\}W$
$=\dfrac{\varepsilon D_{\mathrm{nv}}}{2\hbar}g^{\prime}_{n}\{S_{1}+iS_{2},S_{3}\}+\mathrm{h.c.},$ (40)
where $g^{\prime}_{n}=\left<n\right|\mathrm{exp}\left(-ip_{\xi}\gamma_{0}B/I\omega^{2}\right)\left|n\right$>. For the relevant regime, $g^{\prime}_{n}\simeq 1$, the additional term reduces to
$\delta H_{\mathrm{eff}}=\dfrac{\varepsilon D_{\mathrm{nv}}}{\hbar}\{S_{1},S_{3}\}.$
[…]
$\left|\uparrow^{\prime}\right>=\cos\theta\left|\uparrow\right>+i\sin\theta\left|\downarrow\right>,\quad\left|\downarrow^{\prime}\right>=-\cos\theta\left|\downarrow\right>-i\sin\theta\left|\uparrow\right$>, (41)
with mixing angle $\tan 2\theta=\sqrt{2}\varepsilon D_{\mathrm{nv}}/|\Delta|$. The original basis $\left|\uparrow\right$>, $\left|\downarrow\right$> remains a good approximation whenever
$\theta\simeq\dfrac{\varepsilon D_{\mathrm{nv}}}{\sqrt{2}\Delta}\ll 1\quad\Longleftrightarrow\quad|\Delta|\gg\dfrac{\varepsilon D_{\mathrm{nv}}}{\sqrt{2}}.$ (42)
Numerical simulations (see Fig. 7) show, that for NV-axis deviations $\varepsilon\sim 10^{-2}$, and for magnetic fields $|B|=55\,\mathrm{mT}$ in the dispersive regime, the induced corrections to the spin-rotation dynamics remain negligible on the interferometric timescale.
We note that precise control of NV orientation in nanodiamonds is routinely achievable with current fabrication and pre-selection techniques, ensuring that small misalignments are well within experimental tolerances [30, 78, 79, 80].
D.2 Shape asymmetry
To assess the impact of small deviations from perfect axial symmetry, we extend the model to slightly asymmetric rotors with $I_{2}=I_{1}(1-\delta_{I})$, where $\delta_{I}\ll 1$. This asymmetry adds the kinetic-energy term an additional term $\delta_{I}(\mathbf{J}_{2}-\mathbf{S}_{2})^{2}/2I_{1}$, which weakly couples $\gamma$ to $S_{2}$ and generates a $\gamma$-dependent potential, while the angular momentum component $p_{\alpha}=\mathbf{J}\cdot\mathbf{e}_{z}$ remains conserved. Transforming to the corotating frame and linearizing in $\xi=\pi/2-\beta$, we obtain the correction to Eq. (Gyroscopically stabilized quantum spin rotors),
$\delta H_{\rm rot}$ $=\dfrac{\delta_{I}p_{\xi}^{2}}{2I_{1}}\cos^{2}\gamma+\dfrac{I_{1}\delta_{I}\omega^{2}}{2}(1+\xi^{2})\sin^{2}\gamma-\delta_{I}\omega S_{2}\sin\gamma$
$-\delta_{I}\omega p_{\gamma}\xi\sin^{2}\gamma+\delta_{I}\omega p_{\xi}\sin\gamma\cos\gamma,$ (43)
where we retain only the dominant term coupling $S_{2}$ to the rotational degrees of freedom. The leading effect is the rotation-induced potential $\propto\delta_{I}\omega^{2}\sin^{2}\gamma$, which modifies the $\gamma$ frequency. Importantly, this additional potential does not affect the interferometric protocol, which depends only on the frequency shift induced by the different spin configurations.
For the Barnett-induced alignment along the rotation axis to remain valid, the asymmetry must satisfy
$\delta_{I}\ll\dfrac{\hbar(\omega-\gamma_{0}B)}{I_{1}\omega^{2}}.$ (44)
For single NV centers this condition can be stringent, but in practice the effective Barnett alignment scales with the total magnetic moment. Nanodiamonds hosting multiple NV centers therefore relax this requirement, making the alignment condition easier to satisfy experimentally.
Appendix E Barnett anti-alignment